OpenStudy (anonymous):

Simplify: 33-8 times 3 + 5

OpenStudy (anonymous):

200

OpenStudy (anonymous):

I Think Its 200

OpenStudy (anonymous):

i know its 200

OpenStudy (anonymous):

its multiple choice and the answers you have r a:77 b:3 c:13 d: no answer

OpenStudy (anonymous):

33-8 = 25 3=5 = 8 25 times 8 = 200

OpenStudy (unklerhaukus):

\[33-8 \times 3+5\]\[=33-24+5\]\[=4\] did you forget brackets?

OpenStudy (anonymous):

ok thanks

OpenStudy (unklerhaukus):

=14*